how do you solve this?
-20<(less than or equal to)(I cannot format properly here)4x - 3<(less than or equal to)-2

Found 2 solutions by user_dude2008, RAY100:
Answer by user_dude2008(1862)

-20 <= 4x-3 <= -2

-20+3 <= 4x <= -2+3

-17 <= 4x <= 1

-17/4 <= x <= 1/4 <----- Answer

Answer by RAY100(1637)

-20 <= (4x-3) <= -2
.
add 3 to each area
.
-17 <= 4x<= 1
.
divide by 4
.
-17/4 <= 1x <= 1/4
.
-4.25 <= 1x<= .25
.
.
check,,,x=0,,,-20<=-3<= -2,,,,,,ok
.
x=-4,,,,,-20<=-19<=-2,,,,,,ok
.
x=.1,,,,,,-20<= -2.6 <= -2,,,,,ok
.